Circular No. 5704

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J0422+32 T. Kato, S. Mineshige, and R. Hirata, Kyoto University, write: “V -

band CCD photometry of the optical counterpart of GRO J0422+32 was obtained with the 0.6-m telescope at Ouda Station, Kyoto University, on 14 nights between 1992 Dec. 13 and 1993 Jan. 12. Analysis of the data reveals the appearance of a periodicity other than that reported on IAUC 5676. The present light curve is characterized by roughly sinusoidal variation (amplitude 0.2 mag) and dips. The sinusoidal variation and dips are excellently expressed by a single period 0.2122 ± 0.0001 day. The dips occur exactly at phase 0.44 after the minima of the sinusoidal variation and have typical duration of an hour and an amplitude of 0.1 mag, though varying from 0.05 mag at Jan. 1.52 to 0.17 mag at Jan. 12.58 UT. The firm periodicity and the sharp profile suggest an eclipse origin of the dips. The proposed orbital period (0.2122 day) shows a marked difference from the photometric periodicity of 0.217 day reported on IAUC 5676. The latter is 2.2 ± 0.5 percent longer than the orbital one, thus confirming the superhump nature of the transient periodicity observed in Nov. We suggest that the most probable range of the mass of the compact object is between 2.9 and 6.2 M_{⊙}, following the scheme denoted on IAUC 5676.”

PERIODIC COMET SCHWASSMANN-WACHMANN 1 S. Nakano, Sumoto, Japan, forwards a report from T. Kojima (Chiy-

oda, Gunma Ken) that this comet is again in outburst. A photograph taken with a 0.25-m f/3.4 Wright Schmidt telescope on Jan. 29.53 UT shows the comet at mag 14, being diffuse with strong central condensation, the coma extending to the north in a manner suggesting that the outburst is two or three days old. Other recent total visual magnitude estimates (cf. IAUC 5674, 5675): 1992 Dec. 17.60 UT, 14.1 (A. Nakamura, Kuma, Japan, 0.60m reflector); 25.52, 14.1 (Nakamura); 29.96, 14.5 (H. Mikuz, Ljubljana, Slovenia, 0.20-m reflector + CCD + V filter); 31.00, 14.6 (Mikuz).

NOVA CYGNI 1992 Visual magnitude estimates: 1992 Dec. 23.72 UT, 10.0 (A. Lauvstad,

Levanger, Norway); 1993 Jan. 1.78, 10.0 (P. Schmeer, Bischmisheim, Germany); 15.09, 10.3 (B. H. Granslo, Fjellhamar, Norway); 22.04, 10.5 (W. G. Dillon, Missouri City, TX); Feb. 3.76, 10.5 (Schmeer).

1993 February 4 Daniel W. E. Green